OpenStudy (anonymous):

Assume that the box contains 2 defective widgets and18 acceptable widgets. There are 5 inspectors, each of whom selects a widget from the box at random, inspects it, and then replaces it in the box. Each inspector makes his or her inspection at a different time. What is the probability that at least one inspector will find a defective widget?

OpenStudy (anonymous):

easier to compute the probability that no inspector finds a defective one that should be straightforward

OpenStudy (anonymous):

thats what I thought but my setup is probably not right

OpenStudy (anonymous):

This is my FINAL bernoulli trial problem thank goodness

OpenStudy (anonymous):

probability that any one inspector finds a defective one is \[\frac{1}{9}\] and doesn't find one is \[\frac{8}{9}\] so it should be \[1-(\frac{8}{9})^5\] i think

OpenStudy (anonymous):

"at least one" is a set up for "not none"

OpenStudy (anonymous):

cant get it to work